OpenStudy (anonymous):

Let f(x) = 3x - 6 and g(x) = x - 2. Find f/g and its domain.

OpenStudy (anonymous):

Please explain to me how to get to my answer dont just give it to me

OpenStudy (anonymous):

@agent0smith ?:)

OpenStudy (agent0smith):

f/g means f divided by g. So do that, f is 3x-6, g is x-2. Make sure to divide the entire expression (use parentheses)

OpenStudy (anonymous):

I would divide i just dont know how i would ...like give me an example where you divide?

OpenStudy (agent0smith):

x divided by 2 is: x/2 4x+7 divided by 2x-3 is (4x+7)/(2x-3)

OpenStudy (anonymous):

(3x-6)/(x-2) thats what im trying to solve right? but how? im kinda really dumb

OpenStudy (agent0smith):

Now, get the domain, then we'll simplify. You can't divide by zero, so set x-2 equal to zero and solve for x.

OpenStudy (anonymous):

What does that mean? Omg im so lost

OpenStudy (agent0smith):

x-2 = 0 solve.

OpenStudy (anonymous):

x=2

OpenStudy (agent0smith):

so the domain is all real numbers except x=2 Now, simplify, by factoring out a 2: \[\Large \frac{ 3x-6 }{x-2} = \frac{3(x-2) }{x-2} \]

OpenStudy (anonymous):

I plug in 2 where there is an x?

OpenStudy (agent0smith):

No, that was the domain, we're done with that. Try simplifying it. \[\Large \frac{3(x-2) }{x-2}\]

OpenStudy (anonymous):

Oh okay okay ...umm i would just get left with three right ?

OpenStudy (agent0smith):

Correct. f/g = 3, and the domain is: all real numbers except x=2

OpenStudy (anonymous):

Okay thanks so much :) I swear its great having you fanned lol

OpenStudy (agent0smith):

haha you're welcome :)